# machine learning UWB ultra-wideband

### v1

@@ V-D Fine-tuning WPFMs
Although the foundation model model may already be able to perform general tasks and applications, fine-tuning allows it to adapt the model to specific downstream tasks or domains. To enable this fine-tuning, we believe the following techniques are vital: (i) supervised transfer learning with a small, labeled dataset (tens of samples), (ii) reinforcement learning with human feedback (grading the predictions made by the foundation model before and during model deployment), (iii) zero- and few-shot learning by providing context (data context augmentation), and (iv) example input-output pairs (in-context learning) of the expected downstream task. Additionally, generative AI capabilities can enrich small datasets and allow these be used for fine tuning the model for the considered downstream tasks.
Use case applications With reinforcement learning, a user of an activity recognition system can provide positive or negative feedback, while experts or network controllers can provide efficiency or performance feedback on wireless classification performance. In-context or zero- and few-shot learning can be integrated by giving in and output examples, e.g. a UWB CIR + ’falling down’ movement, as part of a healthcare application. The model can take this example and explain new human activities detected by the UWB radar.
@@ VI Next steps: the road to interactive wireless networks
  • Addressing energy consumption: Foundation models tend to have a large number of parameters, resulting in very high complexity and energy consumption. From a training perspective, trade-offs should be investigated whether this requires more energy than combining multiple traditional task-focused ML solutions. By sharing the WPFM and integrating in multiple domains and applications, the foundation model has the potential to save energy in the end.
  • Embedded and edge-based implementations: Bringing foundation models closer to the end user is a trending topic within the research community. Especially in time critical wireless networks, it is crucial to investigate running the foundation model local on edge or embedded mobile devices, limiting data transfer and latency.

### v2

@@ V-D Fine-tuning WPFMs
Although the foundation model may already be able to perform general tasks and applications, fine-tuning allows it to adapt the model to specific downstream tasks or domains. To enable this fine-tuning, we believe the following techniques are vital: (i) supervised transfer learning with a small, labeled dataset (tens of samples), (ii) reinforcement learning with human feedback (grading the predictions made by the foundation model before and during model deployment), (iii) zero- and few-shot learning by providing context (data context augmentation), and (iv) example input-output pairs (in-context learning) of the expected downstream task. Additionally, generative AI capabilities can enrich small datasets and allow these be used for fine tuning the model for the considered downstream tasks.
Use case applications With reinforcement learning, a user of an activity recognition system can provide positive or negative feedback, while experts or network controllers can provide efficiency or performance feedback on wireless classification performance. In-context or zero- and few-shot learning can be integrated by giving in and output examples, e.g. a UWB CIR + ‘falling down’ movement, as part of a healthcare application. The model can take this example and explain new human activities detected by the UWB radar.
@@ VI Next steps: the road to interactive wireless networks
  • Reasoning capabilities of WPFM and LLMs: To understand why WPFMs and wireless LLMs make certain decisions or descriptions, it becomes crucial to implement reasoning capabilities (such as chain-of-thoughts), which is an active topic of investigation in the generative AI community [sun2023survey].
@@
  • Addressing energy consumption: Foundation models tend to have a large number of parameters, resulting in very high complexity and energy consumption. At inference time, deploying such models in wireless networks is challenging. ML models benefit from reduced wireless bandwidth consumption when they are deployed close to the end device, but these are often resource-constrained [woisetschlager2023federated]. From a training perspective, trade-offs should be investigated whether this requires more energy than combining multiple traditional task-focused ML solutions. By sharing the WPFM and integrating in multiple domains and applications, the foundation model has the potential to save energy in the end.
  • Embedded and edge-based implementations: Bringing foundation models closer to the end user is a trending topic within the research community. In massive (time critical) wireless networks, a large number of data samples are captured, each within their own setting and environment. Implementing federated learning at the edge or embedded devices enables rapid training and reduces the need to send massive amounts of data to a central repository [10078088]. Bringing the knowledge from wireless federated learning into foundation models becomes crucial as the number of unique environmental conditions continues expand.